Net filtration of fluid across the capillaries in muscle is about 0.005 mL min-1 per 100 g of tissue. If glucose is dragged across with this bulk flow, what per cent of glucose transport across the capillary is due to diffusion and what per cent to bulk flow?
Answer #1

Here the relevant equation is

Js = p \DeltaC + JV C

The solute drag term can be calculated as 0.005 mL min-1 /100 g x 90 mg dL-1

The JV term is 5 x 10-5 mL min-1 g-1 and 90 mg dL-1 is equivalent to 4.99 x 10-6 mol mL-1,

so the solvent drag amounts to

5 x 10-5 mL min-1 g-1 x 4.99 x 10-6 mol mL-1 = 24.95 x 10-11 mol min-1g-1

The actual glucose metabolism is given in Part A as 0.0138 x 10-6 mol min-1 g-1 = 1380 x 10-11 mol min-

1 g-1. Thus the fraction delivered by bulk flow is

24.95 x 10-11 mol min-1g-1 / 1380x 10-11 mol min-1 g-1 = 0.0181 = 1.8%